MTH 240 – CH 9 Homework                                                                                                    

For each of these problems follow the 6-step Hypothesis Testing Guideline in the CH 9 Power Point. You can use the p-value or the critical value method. You can calculate the test statistic and p-value (Step 4) either by hand using the formulas and tables, with a graphing calculator, or by using StatCrunch. Include the details of your calculations, not just the values. If you use the critical value method, state the critical value instead of the p-value.

3. Among 13,200 submitted abstracts that were blindly evaluated (with authors and institutions not identified), 26.7% were accepted for publication. Among 13,433 abstracts that were not blindly evaluated, 29.0% were accepted (based on data from “Effect of Blinded Peer Review on Abstract Acceptance,” by Ross, et al., Journal of the American Medical Association, Vol. 295, No. 14). Use a 0.05 significance level to test the claim that the acceptance rate is the same with or without blinding.
Answer #1

Test for P1P2 We want to test whether hypothesis: (Null Hypothesis) (Alternative Hypothesis, also called H1) This is two tail一4.19 Test statistic-4.19 Since this is upper tailed test, p-value = 2*P(Z < |zel) = 2 *P(Z <-4.19) = 0 p-value0 Note that ex